OpenStudy (anonymous):

In a simple random sample of 219 students at a college, 73 reported that they have at least $1000 of credit card debt. What is the 99% confidence interval for the percent of all the students at that college who have at least $1000 in credit card debt?

OpenStudy (amistre64):

what is your idea in creating the interval?

OpenStudy (anonymous):

do i need to find the margin of error?

OpenStudy (amistre64):

margin of error is part of it sure, but in essense: we are trying to find 2, x values, associated with the z scores that canter 99% of the information about the mean. with some adjustments this is just the z formula \[z=\frac{x-\bar x}{\sigma}\] solving for x we get \[\bar x\pm z\sigma=\pm x\] the adjustment is in sigma, we use the standard error instead of the standard devaition

OpenStudy (amistre64):

SE = sqrt(pq/n) for a proportion setup, and xbar = p

OpenStudy (amistre64):

\[p\pm z\sqrt{\frac{pq}{n}}\] this can be stated also as \[\frac {\bar x}{n}\pm z\sqrt{\frac{\bar x~(n-\bar x)}{n^3}}\]

OpenStudy (anonymous):

is it 30.1 and 36.5?

OpenStudy (anonymous):

I'm really confused on this

OpenStudy (amistre64):

weve started with something basic, the z formula, and adapted it to our needs. now all we do is input the information into it. \[\frac {73}{219}\pm z\sqrt{\frac{73~(219-73)}{219^3}}\] and out z value for 99% is what ... 2.567? 2.576? \[\frac {73}{219}\pm 2.576\sqrt{\frac{73~(219-73)}{219^3}}\] multiplying thru by 219 gives us proportions of the sample size needed \[73\pm 2.576\sqrt{\frac{73~(219-73)}{219}}\]

OpenStudy (amistre64):

73 +- 28.8

OpenStudy (amistre64):

and no it cant be 30.1 to 36.5 simply because we are centering around 73 ... 73 is nowhere in between 30 and 36